An electrician wishes to cut a copper wire (ρ=1.724 X10−8Ωm) that has no more than 10Ω of resistance. The wire has a radius of 0.725mm. 

1. The area of cross section of the wire is in square mm


1
Expert's answer
2020-12-15T11:46:51-0500

The area of cross section of the wire is


"A=\pi r^2=3.14\cdot0.725^2=1.65(mm^2)"